OpenStudy (anonymous):

if you know that 7*8=56, how can you use the commutative order property of multiplication to find the product of 8*7?

OpenStudy (inkyvoyd):

uhm.. commutative product states that a*b=b*a... which means that 7*8=8*7=56

OpenStudy (anonymous):

The communtitivue property for multiplication states that the order of the miltiplicands and can be swapped and the resulting product stays the same. So A * B = B * A In your case, that says 7 * 8 = 8 * 7

OpenStudy (anonymous):

Thanks